In a hospital’s shipment of 3 500 insulin syringes, 14 were unusable due to defects. (a) at alpha = 0.05, is this sufficient evidence to reject future shipments from this supplier if the hospital’s quality standard requires 99.7 percent of the syringes to be acceptable?
1
Expert's answer
2021-10-26T08:55:54-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"p\geq0.997"

"p<0.997"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this left-tailed test is "R = \{z: z < -1.6449\}."


The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{3486/3500-0.997}{\sqrt{\dfrac{0.997(1-0.997)}{3500}}}"

"=-1.081747"

Since it is observed that "z = -1.081747 \ge -1.6449=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value is "p =P(Z<-1.081747)= 0.139682," and since "p = 0.139682 \ge 0.05," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than "0.997," at the "\alpha = 0.05" significance level.

It is not sufficient evidence to reject future shipments from this supplier at the "\alpha = 0.05" significance level.